Under normal conditions, the human heart converts about 13.3 J of chemical energy per second into 1.33 W of mechanical power as it pumps blood throughout the body.


(a) Determine the number of Calories required to power the heart for one day, given that 1 Calorie equals 4186 J.

 __Cal


(b) Metabolizing 1 kg of fat can release about 9000 Calories of energy. What mass of metabolized fat (in kg) would power the heart for one day?

__kg


1
Expert's answer
2021-09-27T15:44:17-0400

(a) Human convert = 13.3 J energy per second

In one day "= 13.3 \times 86400 = 1149120 \;J"

"= \frac{1149120}{4186} = 274.52 \;cal"

(b) Proportion:

1 kg = 9000 cal

x - 274.52 cal

"x = \frac{274.52 \times 1 }{9000} = 0.0305 \;kg"